Circular No. 8807

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V1280 SCORPII This nova (cf. IAUC 8803) has brightened greatly in the last week. Se-

lected visual magnitude estimates (some from H. Yamaoka via VSNET, and some sent by E. Waagen, AAVSO): Feb. 8.323 UT, 8.0 (A. Amorim, Florianopolis, Brazil); 10.844, 7.5 (K. Kanai, Gunma, Japan); 11.794, 7.0 (Kanai); 12.326, 6.5 (Amorim); 12.790, 6.2 (Kanai); 12.86, 6.6 (Y. Sakurai, Mito, Ibaraki-ken, Japan; via S. Nakano); 13.246, 6.0 (J. Ripero, Madrid, Spain); 14.782, 5.4 (Kanai); 15.215, 5.2 (Ripero); 15.851, 4.5 (Kanai); 16.212, 4.2 (W. Kriebel, Schierling/Walkenstetten, Germany); 16.323, 3.9 (Amorim); 16.231, 3.8 (P. Schmeer, Bischmisheim, Germany); 16.458, 3.7 (M. Simonsen, Imlay City, MI).

H. Yamaoka, Kyushu University, writes that low-resolution spectroscopy obtained by M. Fujii (Ibara, Okayama, Japan) on Feb. 12.88 and 14.83 UT, and by H. Naito (Nishi-Harima Astronomical Observatory) on Feb. 14.86, shows a much bluer continuum than in the spectroscopy taken on Feb. 5.87 (cf. IAUC 8803); the new spectra resemble the spectrum of an F-type star. Balmer lines (relatively weaker than in the Feb. 5 spectrum) show clear P-Cyg profiles (expansion velocity ∼ 500 km/s), along with other weak lines.

SUPERNOVAE 2006tg–2006ue Twenty supernova discoveries (designated 2006tg–2006tz), all peaking

in the magnitude range R = 20.5–23.5, have been reported by P. Challis on behalf of the ESSENCE team; three were classified as of type II, one of type Ic, and the rest of type Ia (details on CBET 830). Also, five new type-Ia supernovae (designated 2006ua–2006ue) found at Steward Observatory, all in the magnitude range g = 21.5-22.7, were reported by D. Sand et al. (details on CBET 831).

COMET C/2003 C6 (SOHO) An additional Kreutz sungrazing comet has been found on archival

SOHO website images (cf. IAUC 8805). K. Battams notes that C/2003 C6 was stellar in appearance and faint (mag ∼ 8).

Comet 2003 UT α_{2000} δ_{2000} Inst. F MPEC
C/2003 C6 Feb. 8.179 21^{h}41.^{m}0 −16^{o}21′ C3 RK 2007-B75
2007 February 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green